# Blueberry galaxies up to 200 Mpc and their optical/infrared properties

### v1

@@ 3.7 Infrared colors of blueberry galaxies
The WISE IR colors of the aforementioned samples are shown in Figure 9. This comparison involves 38 BBs of this work. Additionally, 39, 719, and 27 galaxies of the Cardamone et al. (2009), Izotov et al. (2011), and Yang et al. (2017) samples respectively that were detected with in WISE bands-1, -2, and -3 in the ALLWISE catalog (Wright et al. 2010; Mainzer et al. 2011) are also examined. The cross-matching with ALLWISE was performed with a search radius of . Additionally, all HECATE galaxies detected with in the same WISE bands, and classified as star-forming by MPA-JHU, are also shown representing the main population of SFGs.
Most BDGs, GPs, and BBs deviate significantly compared to the main population of SFGs, having significantly redder colors both in and . The BBs of this work have , and . Similarly, the Yang et al. (2017) BBs were found to have , and . The majority of all BBs have and . Compared to other dwarf SFGs, BBs show the reddest IR colors, as red as the much dustier and more massive, (ultra) luminous IR galaxies (U-LIRGS; Jarrett et al. 2011). The comparison between the BBs with extended photometry (17 sources with S/N) and compact BBs (17 sources with S/N) shows that the latter have slightly redder IR colors indicating more intense activity (Table 4). However, their IR color distributions overlap significantly.
The BDGs of the larger Izotov et al. (2011) sample are spread over a broad area of the diagram with some reaching extremely red IR color (e.g., ) with most, however, placed in-between the BBs and the typical SFGs. The Cardamone et al. (2009) GPs have similar colors to BBs, however, they show relatively bluer colors.
The dependence of the IR colors of BBs and BDGs on their star-forming intensity is examined in Figure 10. This comparison involves 28, 19, 576, and 27 galaxies of this work, the Cardamone et al. (2009), the Izotov et al. (2011), and the Yang et al. (2017) samples respectively with in the WISE bands-1, -2, and -3. Here, the sSFR is as reported in MPA-JHU for the Cardamone et al. (2009), and Izotov et al. (2011) samples, while for the BBs of this work, the sSFR estimations come from the SED fitting (Section 3.3). The sSFRs of the Yang et al. (2017) BBs are as reported in the same work.
As the color coding reveals, BDGs with higher sSFR tend to have redder IR colors. The population of BDGs shows a stronger color gradient with the color rather than with the color. The majority of BDGs having are galaxies with increased star-forming intensity (log ), however, there are a few high-sSFR, but low-mass BBs, that have .
@@ 4.1 Classification of the Blueberry galaxies
The optical emission-line diagram (BPT; Baldwin et al. 1981; Veilleux & Osterbrock 1987; Kewley et al. 2001; Kauffmann et al. 2003) analysis of this sample classifies all the proposed BBs as purely star-forming galaxies. Similarly, the 4D classification scheme of Stampoulis et al. (2019), which takes into account simultaneously all four emission-line ratios of the traditional BPT diagram, classifies as star-forming with probability all 21 spectroscopically classifiable BBs.
@@
Additionally, to characterize the activity of all the sources of this sample, we apply the scheme proposed by Daoutis et al. (2023) which is based on their photometric IR , , and the visible colors. This method also returns the probability of each source belonging to any of the classes. Again, based on this independent method, all 48 BBs are classified as SFGs. However, only one galaxy, BB 5, was classified with probability to be a SFG which is considered as a secure estimation, while the rest have . Galaxy BB 31/PGC 3108006 has the highest probability of being classified as AGN with but is still considered as a SFG with a much higher probability ().
@@
Although most of the GPs and BBs are above the Stern et al. (2012) color limit all these galaxies are classified as purely star-forming based on the BPT diagram and the IR diagnostics of Daoutis et al. (2023). Additionally, a rather small fraction of the Izotov et al. (2011) BDGs and a larger fraction of the Cardamone et al. (2009) GPs are placed inside the Jarrett et al. (2011) and Satyapal et al. (2018) wedges of AGN galaxies. The Satyapal et al. (2018) criteria were particularly defined to separate the IR colors of BDGs and AGN. To do so, they modeled with Starburst99 BDGs with a continuous starburst at 5 Myrs and a very strong radiation field, and then modeled the nebular and dust emission with Cloudy (Ferland et al. 2013). However still, a small fraction of IR-detected Izotov et al. (2011) BDGs and the majority of Cardamone et al. (2009) GPs fulfill these color criteria, possibly due to complex SFHs and the presence of both old and young SPs whose combined emission can soften the color.
@@
Harish et al. (2023) studied the IR variability of the BPT-star-forming GPs falling within the Jarrett et al. (2011) criteria. Two sources showed variability in their IR emission which is a strong indication of the presence of an AGN. Out of all BBs studied here, only BB 2/PGC 5094058 falls within the Jarrett et al. (2011) wedge, close to its upper limit. Overall, the majority of GPs, and BBs are above the Stern et al. (2012) limit, but most Izotov et al. (2011) BDGs are below. Only a fraction of the BDGs, GPs, and BBs enter the Mateos et al. (2012) wedge. Moreover, while the Jarrett et al. (2011) and Satyapal et al. (2018) color criteria work well for the majority of BBs, they tend to misclassify GP galaxies.
@@
Blueberries BB 22 and BB 7 were reported as AGN galaxies by Shirazi & Brinchmann (2012). These sources were most likely classified as AGN because they show slightly broader emission lines. However, visual inspection of their spectra shows that both permitted and forbidden emission lines show broad features which is a characteristic of outflows related to star-forming activity (e.g., supernovae, stellar winds, superbubbles) rather than an AGN. Using our emission-line flux measurements (Section 3.6), we plot in Figure 11 the – diagram for the BBs of this work. BB 22 is the only source located in the AGN region of the diagram while the rest of BBs are clearly below the Shirazi & Brinchmann (2012) relation.
@@
In order to further examine the classification of the proposed BBs, we performed an archival search of their radio footprints. While the majority of the sources were not observed by LOFAR DR2 (White et al. 1997; Shimwell et al. 2022), which covers part (27%) of the northern sky in low 120–168 MHz frequencies, it successfully detected BBs 7, 8, 9, 21, 27, and 33 but BB 22 was not observed. Among all BBs identified by this work, BB 22 is the only source with detectable radio emission in the radio surveys NRAO VLA Sky Survey (NVSS; Condon et al. 1998), Very Large Array Sky Survey (VLASS; Gordon et al. 2021), and the Rapid ASKAP Continuum Survey (RACS; Hale et al. 2021) that operate in higher radio frequencies. All other BBs were not detected or were not observed from the aforementioned radio surveys even though some should have been detected based on the general SFR–radio relation. However, it has been shown that low-mass galaxies do not follow closely the general SFR-radio relations and are usually underluminous (e.g., Kouroumpatzakis et al. 2021; Borkar et al. 2024). Thus, the radio-detected BB 22 is an exception.
@@
In case BB 22 hosts an AGN we estimate the mass of its central massive black hole (BH). Assuming the typical BH-mass to stellar-mass ratio found in the local Universe (–; e.g., Marconi & Hunt 2003; Kormendy & Ho 2013), the BH mass of BB 22 is estimated to be around – which would place it near the range of intermediate-mass black holes. However, the velocity dispersion of BB 22 was found to be , and based on the M– relation (Tremaine et al. 2002) the BH mass is estimated to be .
@@ 4.2 Stellar populations and interstellar medium of Blueberry galaxies
The majority of BBs have lower metallicities compared to the expected based on their and SFR (Figure 7). A possible explanation could be the very rapid growth of their overall low stellar mass, driven by strong and recent star-formation bursts. Considering that BBs are the most highly star-forming dwarf galaxies in the local Universe, having also very low stellar masses (; Figure 4), a recent star-formation burst could generate a relatively large fraction of their total number of stars. Having very recent and strong starbursts could explain the very low metalicities of BBs, considering the time needed for metals to be distributed in the ISM after the star-formation episodes.
@@
Additionally, most of this work’s BBs show stronger extinction compared to the expected by their metallicity and are overall below the average relation (Figure 8). This result also suggests that these galaxies had very recent starburst episodes. Recent and ongoing star-formation events produce dust. However, because these are very young galaxies with very low stellar mass, dominated by a recent star-formation episode, there were not many cycles of star formation that would allow the dispersion of metals in the ISM. Moreover, these sources must have had very recent star-formation episodes since more than a few Myrs in such strong ionizing radiation fields would lead to the destruction of the dust (e.g., Nozawa et al. 2006; Murata et al. 2014).
@@
The analysis of the BB’s IR photometry is complimentary to this interpretation. The red IR colors of these BBs could be explained based on the galaxies having very recent star-formation episodes. The dominant SPs in BBs are young and massive stars and their strong UV emission is absorbed by the dust which is present in the birth clouds and the ISM, leading to strong emission in the mid-IR part of the spectrum. In contrast, the near-IR which traces the thermal emission of low-mass/older stars is relatively suppressed due to the absence of a strong population of older stars in BBs. Because BBs are galaxies with very low metallicities, we do not expect a relatively high fraction of small dust grains and polycyclic aromatic hydrocarbons (PAHs) to be present (e.g., Draine et al. 2007; Engelbracht et al. 2008; see also the IR spectrum of the low-metallicity BDG UGCA 219 by Brown et al. (2014)), which could otherwise contribute to the mid-IR emission. Thus, the progressively redder IR colors of BBs can be mainly attributed to dust that is stochastically heated by the strong radiation field of their young SPs.
@@
It is known that the spectra of dwarf SFGs tend to have increased intensity in the mid-IR wavelengths (e.g., Brown et al. 2014) while some BDGs can even mimic the IR colors of AGN galaxies (e.g., Satyapal et al. 2014; Hainline et al. 2016; Daoutis et al. 2023). The sources that are placed in the Jarrett et al. (2011), and Satyapal et al. (2018) wedges are mainly BDGs with very high sSFR which are also among the most massive BDGs with similar levels of sSFR. This indicates that the presence of a relatively stronger population of older stars, in combination with the ongoing star formation, could be the reason for the softer that leads BDGs to mimic the IR colors of AGN. The fact that the sSFR of the BDGs correlates more strongly with rather than with (Figure 10) indicates that the intensity of the radiation field may also affect the overall shape of their IR spectrum. In contrast to the more massive BDGs, the very-low mass BBs tend to have a more stochastic behavior but most of them are placed in the upper right part of the diagram and do not enter the Jarrett et al. (2011), and Satyapal et al. (2018) wedges.
4.3 Comparison between Blueberry and Green Pea galaxies.
@@
Blueberry galaxies are considered the lower-mass and nearby analogs of GPs. Although generally, BBs are smaller and less massive, they also show lower metallicities compared to GPs (Figure 7) and have on average about an order of magnitude higher sSFR (Figures 4, 6). Moreover, the BBs’ higher values of EWs in all the compared emission lines indicate that they are forming stars even faster and their underlying population of older stars is smaller. It has been shown that there is a tight correlation between the and sSFR of galaxies (e.g., Belfiore et al. 2018). The EW takes into account the emission line component over the continuum at the specific wavelength of each line. As the continuum traces the stellar emission of low-mass SPs, and the emission lines trace the ionizing radiation of the massive stars, the EW is also a tracer of the relative star-forming activity over the stellar component of galaxies.
@@
The comparison of the samples’ spectroscopic properties reveals that on average BBs show stronger levels of ionization (Figure 6). Additionally, most of the spectroscopically observed BBs of this work show in emission, a high ionization emission line that, in the absence of AGN, requires the presence of many very massive and young stars. The comparison of the IR colors of the two populations is complementary to this conclusion since BBs show on average redder IR colors, mainly in the color (Figure 9).
@@
Although the GPs selection process followed by Cardamone et al. (2009) was not restrictive on how blue the galaxies’ colors could be, their selected sources were limited to the redshift range . This led to the exclusion of BB-like sources because BBs, although bluer, are less luminous compared to the more massive GPs. For instance, the brightest source of this sample, BB 22 (with observed mag), would be marginally detectable ( mag), and the second brightest source, BB 32 (with observed mag), would be completely undetectable ( mag) by SDSS if they were placed at Mpc (z=0.3), a distance representative of most GPs. Thus, although BBs and GPs are part of the larger population of BDGs, the particular samples have intrinsic differences with the first being smaller and overall more extreme.

### v3

@@ 3.7 Infrared colors of blueberry galaxies
The WISE IR colors of the aforementioned samples are shown in Figure 9. The comparison shown in the figure involves 38 BBs of this work. Also included are the 39, 719, and 27 galaxies respectively of the Cardamone et al. (2009), Izotov et al. (2011), and Yang et al. (2017) samples that were detected with an in WISE bands 1, 2, and 3 in the ALLWISE catalog (Wright et al. 2010; Mainzer et al. 2011). We performed cross-matching with ALLWISE with a search radius of . Additionally, all HECATE galaxies detected with an in the same WISE bands and classified as star forming by MPA-JHU are also shown, representing the main population of SFGs.
Most BDGs, GPs, and BBs deviate significantly compared to the main population of SFGs, having significantly redder colors both in and . The BBs of this work have and . Similarly, the Yang et al. (2017) BBs were found to have and . The majority of all the BBs have and . Compared to other dwarf SFGs, BBs show the reddest IR colors and are as red as the much dustier and more massive (ultra) luminous IR galaxies (U-LIRGS; Jarrett et al. 2011). A comparison between the BBs with extended photometry (17 sources with S/N) and the compact BBs (17 sources with S/N) showed that the latter have slightly redder IR colors, indicating more intense activity (Table 4). However, their IR color distributions overlap significantly.
The BDGs of the larger Izotov et al. (2011) sample are spread over a broad area of the diagram, with some reaching an extremely red IR color (e.g., ), but most are between the BBs and the typical SFGs. The Cardamone et al. (2009) GPs have colors similar to BBs; however, they show relatively bluer colors.
The dependence of the IR colors of BBs and BDGs on their star-forming intensity is examined in Figure 10. This figure presents a comparison that involves 28 galaxies of this work and 19, 576, and 27 galaxies respectively of the Cardamone et al. (2009), Izotov et al. (2011), and Yang et al. (2017) samples with an in the WISE bands 1, 2, and 3. Here, the sSFR is as reported in MPA-JHU for the Cardamone et al. (2009) and Izotov et al. (2011) samples, while for the BBs of this work, the sSFR estimations come from the SED fitting (Section 3.3). The sSFRs of the Yang et al. (2017) BBs are as reported in that work.
As the color coding reveals, BDGs with a higher sSFR tend to have redder IR colors. The population of BDGs shows a stronger color gradient with the color than with the color. Most of the BDGs with are galaxies with increased star-forming intensity (log ). However, there are a few high-sSFR but low-mass BBs that have .
@@ 4.1 Classification of the blueberry galaxies
The optical emission-line diagram (BPT; Baldwin et al. 1981; Veilleux & Osterbrock 1987; Kewley et al. 2001; Kauffmann et al. 2003) analysis of this sample classified all the proposed BBs as purely SFGs. Similarly, the 4D classification scheme of Stampoulis et al. (2019), which simultaneously takes into account all four emission-line ratios of the traditional BPT diagram classified as star forming with probability all 21 spectroscopically classifiable BBs.
@@
Additionally, to characterize the activity of all the sources of this sample, we applied the scheme proposed by Daoutis et al. (2023), which is based on their photometric IR , and the visible colors. This method also returns the probability of each source belonging to any of the classes. Again, based on this independent method, all 48 BBs were classified as SFGs. However, only one galaxy, BB 5, was classified as an SFG with , which is considered as secure estimation, while the rest have . Galaxy BB 31 (PGC 3108006) has the highest probability of being classified as an AGN, with , but it is still considered an SFG with a much higher probability ().
@@
Although most of the GPs and BBs are above the Stern et al. (2012) color limit, all of these galaxies were classified as purely star forming based on the BPT diagram and the IR diagnostics of Daoutis et al. (2023). Additionally, a rather small fraction of the Izotov et al. (2011) BDGs and a larger fraction of the Cardamone et al. (2009) GPs were placed inside the Jarrett et al. (2011) and Satyapal et al. (2018) wedges of AGN galaxies. The criteria of Satyapal et al. (2018) were particularly defined to separate the IR colors of BDGs and AGNs. To do so, they modeled with Starburst99 BDGs with a continuous starburst at 5 Myrs and a very strong radiation field, and then they modeled the nebular and dust emission with Cloudy (Ferland et al. 2013). However, a small fraction of IR-detected Izotov et al. (2011) BDGs and the majority of Cardamone et al. (2009) GPs still fulfilled these color criteria, possibly due to complex SFHs and the presence of both old and young SPs, whose combined emission can soften the color.
@@
Harish et al. (2023) studied the IR variability of the BPT star-forming GPs falling within the Jarrett et al. (2011) criteria. Two sources showed variability in their IR emission, which is a strong indication of the presence of an AGN. Out of all BBs studied here, only BB 2 (PGC 5094058) falls within the Jarrett et al. (2011) wedge, close to its upper limit. Overall, the majority of GPs and BBs are above the Stern et al. (2012) limit, but most Izotov et al. (2011) BDGs are below it. Only a fraction of the BDGs, GPs, and BBs enter the Mateos et al. (2012) wedge. Moreover, while the Jarrett et al. (2011) and Satyapal et al. (2018) color criteria work well for the majority of BBs, they tend to misclassify GP galaxies.
@@
The galaxies BB 22 and BB 7 were reported as AGN galaxies by Shirazi & Brinchmann (2012). These sources were most likely classified as AGNs because they show slightly broader emission lines. However, visual inspection of their spectra shows that both permitted and forbidden emission lines show broad features, which is a characteristic of outflows related to star-forming activity (e.g., supernovae, stellar winds, superbubbles) rather than an AGN. Using our emission-line flux measurements (Section 3.6), we plot in Figure 11 the – diagram for the BBs of this work. Notably, BB 22 is the only source located in the AGN region of the diagram, while the rest of BBs are clearly below the Shirazi & Brinchmann (2012) relation.
@@
In order to further examine the classification of the proposed BBs, we performed an archival search of their radio footprints. While a majority of the sources were not observed by LOFAR DR2 (White et al. 1997; Shimwell et al. 2022), which covers part (27%) of the northern sky in low 120–168 MHz frequencies, it successfully detected BBs 7, 8, 9, 21, 27, and 33, but BB 22 was not observed. Among all BBs identified by this work, BB 22 is the only source with detectable radio emission in the radio surveys of the NRAO VLA Sky Survey (NVSS; Condon et al. 1998), Very Large Array Sky Survey (VLASS; Gordon et al. 2021), and the Rapid ASKAP Continuum Survey (RACS; Hale et al. 2021) that operate at higher radio frequencies. All other BBs were not detected or were not observed by the aforementioned radio surveys even though some should have been detected based on the general SFR–radio relation. However, it has been shown that low-mass galaxies do not closely follow the general SFR-radio relations and are usually underluminous (e.g., Kouroumpatzakis et al. 2021; Borkar et al. 2024). Thus, the radio-detected BB 22 is an exception.
@@
In case BB 22 hosts an AGN, we estimated the mass of its central massive black hole (BH). Assuming the typical BH mass to stellar mass ratio found in the local Universe (–; e.g., Marconi & Hunt 2003; Kormendy & Ho 2013), the BH mass of BB 22 is estimated to be around –, which would place it near the range of intermediate-mass BHs. However, the velocity dispersion of BB 22 was found to be , and based on the M– relation (Tremaine et al. 2002), the BH mass is estimated to be .
@@ 4.2 Stellar populations and interstellar medium of blueberry galaxies
The majority of BBs have lower metallicities compared to the expected values based on their and SFR (Figure 7). A possible explanation for this could be the very rapid growth of their overall low stellar mass, driven by strong and recent star-formation bursts. Considering that BBs are the most highly star-forming dwarf galaxies in the local Universe and also have very low stellar masses (; Figure 4), a recent star-formation burst could generate a relatively large fraction of their total number of stars. Having very recent and strong starbursts could explain the very low metallicities of BBs, considering the time needed for metals to be distributed in the ISM after the star-formation episodes.
@@
Additionally, most of this work’s BBs show a stronger extinction compared to what is expected based on their metallicity and are overall below the average relation (Figure 8). This result also suggests that these galaxies had very recent starburst episodes. Recent and ongoing star-formation events produce dust. However, because these galaxies are very young and have a very low stellar mass, dominated by a recent star-formation episode, there have not been many cycles of star formation that would have allowed for the dispersion of metals in the ISM. Moreover, these sources must have had very recent star-formation episodes since more than a few million years passing in such strong ionizing radiation fields would lead to the destruction of the dust (e.g., Nozawa et al. 2006; Murata et al. 2014).
@@
The analysis of the BB’s IR photometry is complementary to this interpretation. The red IR colors of these BBs could be explained based on the galaxies having very recent star-formation episodes. The dominant SPs in BBs are young and massive stars, and their strong UV emission is absorbed by the dust that is present in the birth clouds and the ISM, leading to strong emission in the mid-IR part of the spectrum. In contrast, the near-IR, which traces the thermal emission of low-mass and long-lived stars, is relatively suppressed due to the absence of a strong population of older stars in BBs. Because BBs are galaxies with very low metallicities, we did not expect a relatively high fraction of small dust grains and polycyclic aromatic hydrocarbons (PAHs) to be present (e.g., Draine et al. 2007; Engelbracht et al. 2008; see also the IR spectrum of the low-metallicity BDG UGCA 219 by Brown et al. (2014)), which could otherwise contribute to the mid-IR emission. Thus, the progressively redder IR colors of BBs can be mainly attributed to dust that is stochastically heated by the strong radiation field of their young SPs.
@@
It is known that the spectra of dwarf SFGs tend to have increased intensity in the mid-IR wavelengths (e.g., Brown et al. 2014), while some BDGs can even mimic the IR colors of AGN galaxies (e.g., Satyapal et al. 2014; Hainline et al. 2016; Daoutis et al. 2023). The sources that are placed in the Jarrett et al. (2011) and Satyapal et al. (2018) wedges are mainly BDGs with a very high sSFR, and they are also among the most massive BDGs with similar levels of sSFR. This indicates that the presence of a relatively stronger population of older stars, in combination with the ongoing star formation, could be the reason for the softer that leads BDGs to mimic the IR colors of AGNs. The fact that the sSFR of the BDGs correlates more strongly with rather than with (Figure 10) indicates that the intensity of the radiation field may also affect the overall shape of their IR spectrum. In contrast to the more massive BDGs, the very-low-mass BBs tend to have a more stochastic behavior, but most of them are placed in the upper-right part of the diagram and do not enter the Jarrett et al. (2011) and Satyapal et al. (2018) wedges.
4.3 Comparison between blueberry and green pea galaxies
@@
Blueberry galaxies are considered the lower-mass and nearby analogs of GPs. Although BBs are generally smaller and less massive, they also show lower metallicities compared to GPs (Figure 7) and have on average about an order of magnitude higher sSFRs (Figures 4, 6). Moreover, the higher values of EWs of BBs in all the compared emission lines indicate that they are forming stars even faster and that their underlying population of older stars is smaller. It has been shown that there is a tight correlation between the and sSFR of galaxies (e.g., Belfiore et al. 2018). The EW takes into account the emission-line component over the continuum at the specific wavelength of each line. As the continuum traces the stellar emission of low-mass SPs and the emission lines trace the ionizing radiation of the massive stars, the EW is also a tracer of the relative star-forming activity over the stellar component of galaxies.
@@
Comparison of the samples’ spectroscopic properties reveals that, on average, BBs show stronger levels of ionization (Figure 6). Additionally, most of the spectroscopically observed BBs of this work show in emission, a high ionization emission line that, in the absence of an AGN, requires the presence of many very massive and young stars. The comparison of the IR colors of the two populations is complementary to this conclusion since BBs show on average redder IR colors, mainly in the color (Figure 9).
@@
Although the GP selection process followed by Cardamone et al. (2009) was not restrictive regarding how blue the galaxies’ colors could be, their selected sources were limited to the redshift range . This led to the exclusion of BB-like sources because BBs, although bluer, are less luminous compared to the more massive GPs. For instance, the brightest source of this sample, BB 22 (with observed mag), would be marginally detectable ( mag), and the second brightest source, BB 32 (with observed mag), would be completely undetectable ( mag) by SDSS if they were placed at Mpc (z=0.3), a distance representative of most GPs. Thus, although BBs and GPs are part of the larger population of BDGs, the particular samples have intrinsic differences, with the former being smaller and overall more extreme.